### v3

# A longitudinal Bayesian framework for estimating causal dose-response relationships

Abstract
Existing causal methods for time-varying exposure and time-varying confounding focus on estimating the average causal effect of a time-varying binary treatment on an end-of-study outcome, offering limited tools for characterizing marginal causal dose–response relationships under continuous exposures. We propose a scalable, nonparametric Bayesian framework for estimating marginal longitudinal causal dose–response functions with repeated outcome measurements. Our approach targets the average potential outcome at any fixed dose level and accommodates time-varying confounding through the generalized propensity score. The proposed approach embeds a Dirichlet process specification within a generalized estimating equations structure, capturing temporal correlation while making minimal assumptions about the functional form of the continuous exposure. We apply the proposed methods to monthly metro ridership and COVID-19 case data from major international cities, identifying causal relationships and the dose–response patterns between higher ridership and increased case counts.
Keywords: Dose-response relationship; Bayesian bootstrap; Dirichlet processes; Longitudinal data analysis; COVID-19; Transportation engineering
1 Introduction
The SARS-CoV-2 virus pathogen, which causes the COVID-19 disease, is considered to transmit via two main pathways: (i) directly through droplets and aerosols, and (ii) indirectly through fomite transmission (for example, Derqui et al., 2023). Public transport modes serving dense urban areas induce prolonged exposure for passengers in enclosed and oftentimes poorly ventilated spaces, and they also facilitate the mixing of populations at the origin and destination trip-ends. Consequently, we can reasonably hypothesize that public transport is a vector of virus transmission in a city. While this hypothesis seems intuitively reasonable, and has been put forward previously, including in the mass media (Thomas et al., 2022), it remains largely unsubstantiated statistically under the causal lens. Existing empirical studies quantifying the dynamic impact of mass public transport on COVID-19 transmission are scarce. Most previous studies have analyzed aggregate human mobility, rather than mass transit ridership specifically, for instance, using less granular mobile phone data sources such as those available via Google Mobility or Apple Maps. Despite growing evidence of multi-layered causal pathways, the causal dynamics connecting mass transit ridership, such as metro ridership, to infectious-disease spread remain underexplored.
Using the standard causal inference notation, we define our problem in relation to the triple $(Y,D,X)$, indexed by $t$ to denote time: where $Y$ is the outcome (COVID-19 cases in the city), $D$ is the treatment (urban mass transit ridership), $X$ is a set of time-varying confounders and $U$ is a set of confounders that are time-invariant. We are interested in the causal effects of $D$ on $Y$. Therefore, we face two key challenges in estimating this marginal causal effect. First, we must adjust for the existence of confounding influences, $X$ and $U$, which determine both transit ridership and disease propagation. These include time-varying (i.e., $X$) factors such as the severity of containment measures on human mobility (e.g. lockdown measures), urban mobility characteristics, and vaccination rates, along with a set of unmeasured time-invariant (i.e. $U$) city level attributes such as characteristics of the built-environment, urban culture, and baseline socio-demographics. Second, there are dynamic effects presented such that past outcomes act as contemporaneous confounders. Thus, we observe a repeatedly measured outcome, a time-dependent treatment (dose) and a set of dynamic time-varying confounders.
Most existing methods for causal estimation of the effect of a continuous-exposure are developed for the point-treatment framework, including the generalized propensity score (GPS)-based approaches (Robins et al., 2000; Hirano and Imbens, 2004; Moodie and Stephens, 2012; Galvao and Wang, 2015), g-computation and outcome regression-based approaches (Hill, 2011), and a doubly robust estimator considering combining both treatment and outcome models (Kennedy et al., 2017). Extensions to longitudinal causal data are very limited and restricted to the parametric frequentist paradigms (Robins et al., 2000; Vansteelandt and Sjolander, 2016). In particular, g-computation can be computationally intensive and intractable when estimating treatment regime effects in longitudinal settings with high-dimensional time-varying confounding, as it requires fitting sequential models of the complete data-generating process and simulating counterfactual trajectories of time-varying covariates and outcomes at a set of treatment levels (Robins, 1986; Linero, 2022). On the other hand, the Bayesian lens for longitudinal causal dose-response analysis is largely missing and remains underexplored in the literature. The Bayesian estimation framework can offer both statistical and practical advantages, including the propagation of uncertainty through complex model structures and the ability to provide probabilistic summaries of causal effects for intuitive causal decision-making in applied research, which facilitates decision-making in applied settings such as our ridership–COVID study. In addition, frequentist methods such as generalized estimating equations (GEEs) do not naturally accommodate prior information. These features motivate a Bayesian alternative that retains the scalability of estimating-equation methods while enabling posterior summaries of longitudinal dose–response effects. The incorporation of prior beliefs under the proposed Bayesian approach can be particularly valuable in biomedical and policy applications where historical evidence or expert knowledge is available. Moreover, unlike approaches that rely on large-sample approximations, Bayesian inference provides coherent uncertainty quantification even in small-sample settings, which commonly arise in dose–response analyses. These considerations motivate us to develop Bayesian alternatives for longitudinal causal dose–response modeling. In this paper, therefore, we fill this literature gap by introducing a flexible yet computationally tractable non-parametric Bayesian estimation framework that estimates the entire longitudinal dose-response surface while making minimal distribution assumptions about the functional form of the continuous exposure.
Our specific methodological contributions in this paper are on three fronts: i) we propose a nonparametric Bayesian framework using the Dirichlet process (DP) for marginal causal dose–response estimation with longitudinal data, which integrates GEEs with coherent Bayesian uncertainty quantification; ii) we demonstrate how the proposed DP–based framework accommodates multiple causal adjustment strategies, including covariate-adjusted outcome regression and inverse probability weighted marginal structural models; and iii) we provide computationally scalable algorithms and open-source code (https://github.com/yumcgill/Bayes_DoseResponse) making the proposed Bayesian approach practical for large, real-world datasets for wide adoption and application beyond the motivating study.
The remainder of this paper is organized as follows. In Section 2, we review methods for causal analysis of longitudinal data with repeated outcomes and introduce a non-parametric Bayesian formulation for inference when the model is specified through GEEs. Section 3 presents simulation studies, comparing a weighting-based method with an approach that incorporates the GPS as a covariate. In Section 4, we apply our methods to a real data example, analyzing the causal effect of mass transit ridership on the spread of COVID-19. Finally, we conclude with some remarks in Section 5.
2 Methodology
2.1 Causal inference for marginal dose–response functions with repeated outcomes
In causal inference for the longitudinal setting, for $t=1,\ldots,T$, let $D_{t}\in\mathbb{R}$ denote the (possibly continuous) dose at time $t$, Let $X_{t}$ denote a set of time-varying observed covariates, and let $U$ denote an unobserved time-invariant baseline covariate that may affect both treatment assignment and outcomes. Therefore, the observed data consist of the longitudinal process $Z=\{(Y_{t},X_{t},D_{t}):t=1,\ldots,T\}$. Using the potential outcome framework for longitudinal data, we define $Y_{t}(d_{t})$ as the counterfactual outcome at time $t$ that would be observed if a study unit were exposed to the dose $d_{t}$ at time $t$. Our primary estimand is a marginal dose–response function, defined as
$\mu(d)=\mathbb{E}\!\left\{Y(d)\right\},$
where $Y(d)$ denotes the potential outcome that would be observed if, at each time point, the exposure were set to the constant level $d$. This estimand represents the average potential outcome (APO) under an exposure level $d$. In other words, the estimand depends only on the marginal level of exposure $d$, rather than on the full longitudinal treatment path used to achieve it. This marginal dose–response parameterization was commonly used in transportation engineering. As in the transportation setting, it is neither feasible nor substantively meaningful to define a single, well-specified longitudinal treatment regime. Instead, our scientific goal is to quantify the APO at a given ridership on COVID-related cases. Therefore, we believe it is suitable to estimate the dose-response relationship in our COVID analysis. For example it was employed by Graham et al. (2014) in their analysis to quantify the effect of the network capacity expansions on urban traffic volumes. Therefore, we adopt this estimand to characterize the dose–response relationship in our COVID analysis. In Appendix B, an extension of this framework is provided for estimating regime-specific APO associated with identification results.
To identify the APO $\mu(d)=\mathbb{E}\{Y(d)\}$, we invoke the following assumptions. First, we assume no interference across units, i.e., for each study unit $i$ ($i=1,\ldots,n$), $Y_{i}(d_{1},\ldots,d_{n})=Y_{i}(d_{i})$, corresponding to the stable unit treatment value assumption (SUTVA). Second, we assume consistency, that is, if an study unit is observed to receive dose $D_{t}=d$ at time $t$, then the observed outcome satisfies $Y_{t}=Y_{t}(d)$. Third, we assume marginal ignorability of treatment assignment: at each time $t=1,\ldots,T$, $Y(d)\;\perp\!\!\!\perp\;D_{t}\mid X_{t}$. This assumption does not condition on past dose, and allows past dose to affect both $D$ and $Y$ indirectly, provided those effects are captured by $X$. Fourth, we assume positivity. For any covariate value $x_{t}$ with positive density, $f(d_{t}\mid X_{t}=x_{t})>0$ over the range of doses considered. Finally, we assume that unobserved time-invariant heterogeneity $U$ is mean-independent of the observed time-varying covariates, i.e., $\mathbb{E}(U\mid X_{t})=0$. Figure 1 illustrates a general longitudinal causal structure consistent with these assumptions, allowing for treatment–confounder feedback and unobserved time-invariant heterogeneity and targeting a marginal estimand.
$X_{t}$$D_{t}$$Y_{t}$$U$
Figure 1: Directed acyclic graph for identification of the marginal dose–response function $\mu(d)=\mathbb{E}\{Y(d)\}$ under marginal ignorability. Observed covariates $X_{t}$ block all backdoor paths between exposure $D_{t}$ and outcome $Y_{t}$, while unobserved time-invariant heterogeneity $U$ is orthogonal to $X_{t}$.
Following this marginal effect parameterization, the APO can be modeled as
$g\!\left(\mathbb{E}\!\left\{Y(d)\right\}\right)=m(d;\xi),$
where $g(\cdot)$ is a known link function and $m(\cdot;\xi)$ is a marginal dose–response model indexed by parameters $\xi$. Marginal structural models (MSMs) estimate $\mu(d)$ by reweighting observations using inverse probability weights to remove time-dependent confounding. As an alternative to reweighting, Graham et al. (2014) proposed incorporating the GPS directly into the outcome model. Under correct specification of the GPS model, this approach yields consistent estimation of the marginal dose-response function. Accordingly, the outcome model may be written as
$g\!\left(\mathbb{E}[Y_{t}\mid D_{t},X_{t},\widehat{e}_{t}(X_{t})]\right)=m(D_{t};\xi)+h\!\left(\widehat{e}_{t}(X_{t})\right),$
where $\widehat{e}_{t}(\cdot)$ denotes the estimated GPS at time $t$ and $h(\cdot)$ is a prespecified adjustment function, which may be estimated flexibly (e.g., using splines). By integrating over time and covariate, estimation of $\xi$ yields the time-invariant APO $\mu(d)$. For a continuous exposure, the GPS is defined as $e_{t}(X_{t})=f(D_{t}\mid X_{t})$, where $f(D_{t}\mid X_{t})$ denotes the conditional density of the treatment given covariates. In practice, the GPS is estimated by specifying a marginal mean model for $D_{t}$ given $X_{t}$ and estimating its parameters using the GEE to account for within-subject dependence across repeated treatment measurements. Importantly, although the treatment assignment may depend on an unobserved time-invariant factor $U$, the GEE estimator targets the marginal mean $E(D_{t}\mid X_{t})$, which integrates over the distribution of $U$. Under the assumption that $E(U\mid X_{t})=0$, the resulting estimating equations are unbiased, and the GEE estimator yields consistent estimates of the marginal treatment model parameters regardless of the working correlation structure. The estimated marginal mean and variance are then used to evaluate the conditional density of $D_{t}$ (e.g., assuming normality) at the observed exposure level, yielding a consistent estimate of the GPS. Incorporating this consistently estimated GPS into the outcome GEE therefore yields consistent estimation of the APO. Appendix A provides formal results demonstrating that unobserved time-invariant effects $U$ are integrated out in the marginal treatment model and therefore do not bias the GEE-based estimation of the GPS. Under assumptions, the dose–response function $\mu(d)$ is identified by
$\mathbb{E}[Y_{t}(d)]=\mathbb{E}_{X_{t}}\left[\mathbb{E}\left[Y_{t}(d)\right]\right]=\mathbb{E}_{X_{t}}\left[\mathbb{E}\left[Y_{t}\mid D_{t}=d,e_{t}(d,X_{t})\right]\right],\qquad d\in\mathcal{D},$ (1)
where $e_{t}(d,X_{t})=f(D_{t}=d\mid X_{t})$ and can be consistently estimated via the GEE treatment model. Therefore, the marginal treatment effect associated with parameter $\xi$, can then be estimated using a GEE (Preisser et al., 2002), defined by the following estimating equation
$\sum_{i=1}^{n}\mathbf{U}_{i}(\xi)\equiv\sum_{i=1}^{n}\left(\frac{\partial\mu_{i}(\xi)}{\partial\xi}\right)^{\top}\Sigma_{i}^{-1}\left(Y_{i}-\mu_{i}(\xi)\right)=0,\;\;\;\mu_{i}(\xi)=\mathbb{E}[Y_{t}\mid D_{t},X_{t},\widehat{e}_{t}(X_{t})]$ (2)
where $Y_{i}$}} is a vector of observed response for ithith study unit. This connects the targeting parameter $\xi$ from a link function, and $\Sigma_{i}$ is the $T\times T$ working covariance matrix for study unit $i$. However, as noted in Tchetgen et al. (2012), the GEE approach consistently estimates the parameters of a correctly specified outcome model in absence of weights, irrespective of whether the working correlation structure is correct. The GEE framework offers flexibility in targeting the marginal treatment effect for repeated measurements. However, its reliance on estimating equations, rather than full likelihoods, poses challenges for conventional Bayesian analysis, as there is no explicit distributional form for the marginal mean model. This motivates the development of a fully Bayesian approach, which we outline in the next section.
2.2 Bayesian non-parametric method for GEE-based causal estimands
The Bayesian framework offers calibrated uncertainty quantification and prediction, which can render a more thorough understanding of the causal effect from a practical perspective. There is growing interest in the application of Bayesian methodology to causal problems, but most proposed methods appeal to the principles of PS-adjustment or flexible outcome modeling (Stephens et al., 2023; Li et al., 2023). An advantage of the Bayesian approach for estimation of these models is that estimated causal effects are in the form of distributions, rather than point estimates, allowing us to make probability statements about the causal quantities of interest (i.e., APOs). In fully parametric Bayesian analyses, this is typically achieved by specifying likelihood models for both the outcome and treatment assignment mechanisms. However, such an approach is not directly applicable in the present setting, as the GEE in (2) specifies only the conditional mean structure of the outcome and does not correspond to a fully specified likelihood. Consequently, Bayesian inference in this context must proceed under an approximate (or partially specified) model, which may be misspecified relative to the true data-generating mechanism. Rather than modeling the full joint distribution of the data, inference targets a causal estimand defined through unbiased estimating equations. This perspective aligns with the generalized Bayesian framework and recent work on Bayesian inference for loss functions (Luo et al., 2023). We define the target causal parameter $\xi_{0}$ at the population level through the moment condition
$\mathbb{E}_{F_{0}}\!\left\{\mathbf{U}(\xi_{0})\right\}=0,$ (3)
where $F_{0}$ denotes the true, unknown data-generating distribution and $\mathbf{U}(\cdot)$ is the GEE function defined in (2). Equivalently, $\xi_{0}$ may be characterized as the minimizer of an expected loss function,
$\xi_{0}=\arg\min_{\xi\in\Xi}\;\mathbb{E}_{F_{0}}\!\left\{\ell(Z;\xi)\right\},$ (4)
where $\ell(\cdot)$ is a differentiable loss function whose gradient yields the estimating equations in (3). This loss-based definition is consistent with the generalized Bayesian framework of Luo et al. (2023) and provides a precise definition of the causal estimand without requiring a likelihood. In the Bayesian non-parametric framework, uncertainty about $F_{0}$ is represented through a posterior distribution over probability measures. If we assume that the data points are realizations from a multinomial model on the finite set of the observed data $\left\{z_{1},\ldots,z_{n}\right\}$ with unknown probability $\mathbf{p}=\left(p_{1},\ldots,p_{n}\right)$, and assuming a priori that $\mathbf{p}\sim\text{Dirichlet}\left(\alpha,\ldots,\alpha\right)$, then, a posteriori $\mathbf{p}\sim\text{Dirichlet}\left(\alpha+1,\ldots,\alpha+1\right)$. The standard Bayesian bootstrap is obtained under the improper specification $\alpha=0$ (Rubin, 1981). Therefore, if we repeatedly sample $\mathbf{p}\sim\text{Dirichlet}\left(1,\ldots,1\right)$, then
$F^{(b)}=\sum_{i=1}^{n}p_{i}^{(b)}\delta_{Z_{i}}.$
For each posterior draw $F^{(b)}$, the corresponding parameter draw $\xi^{(b)}$ is defined as the solution to the posterior predictive estimating equation
$\int\mathbf{U}(\xi^{(b)})\,dF^{(b)}(z)=\sum_{i=1}^{n}p_{i}^{(b)}\mathbf{U}_{i}(\xi^{(b)})=0.$ (5)
This construction induces a posterior distribution over $\xi$ through the posterior over $F$, yielding valid Bayesian uncertainty quantification for GEE-defined causal estimands (Newton and Raftery, 1994; Chamberlain and Imbens, 2003). Importantly, the definition of $\xi_{0}$ in (3) and the Bayesian bootstrap implementation in (5) represent equivalent formulations of the same inferential target. The former defines the causal estimand at the population level, while the latter provides a practical posterior sampling scheme. This approach yields a coherent Bayesian non-parametric framework for inference on GEE-based causal parameters, accommodates model misspecification, and avoids the need to specify a full likelihood while retaining the interpretability of marginal causal estimands.
2.3 The generalized Bayesian bootstrap
The previous section described a standard Bayesian bootstrap approach for obtaining a posterior distribution for parameters defined through estimating equations. In our setting, the model is specified through a two–stage system of GEEs, consisting of a GPS model in the first stage and an outcome model in the second stage. Specifically, let $\mathbf{U}_{1}(\cdot)$ denote the GEE associated with the GPS model and $\mathbf{U}_{2}(\cdot)$ denote the GEE for the outcome model. Under the Bayesian bootstrap, we repeatedly sample weights $\mathbf{p}=(p_{1},\ldots,p_{n})\sim\mathrm{Dirichlet}(1,\ldots,1)$, and for each draw of $\mathbf{p}$, solve the weighted estimating equations
$\sum_{j=1}^{n}p_{j}\,\mathbf{U}_{1}(x_{j},d_{j},\gamma)$ $=\mathbf{0},$ (6)
[…]
where $e(x_{j};\gamma)$ denotes the GPS model, and $\widehat{\gamma}$ is the solution to (6). The resulting solution $\widehat{\xi}(\mathbf{p})$ defines a draw from the posterior distribution of $\xi$ induced by the Bayesian bootstrap. Repeating this procedure yields an empirical approximation to the posterior distribution of $\xi$. While this Bayesian bootstrap specification provides a convenient and likelihood-free way to propagate uncertainty through the two-stage GEE system, it also has an important limitation. The Bayesian bootstrap places all its prior mass on the empirical support of the observed data, implicitly assuming that the unknown data-generating distribution $F$ is discrete with atoms fixed only at the observed points. As a result, posterior inference for $\mu(d)$ is entirely driven by reweighting the observed sample and cannot extrapolate beyond the empirical support of the dose. It often tends to underestimate uncertainty in regions where data are sparse, particularly in the tails of the exposure distribution.
As proposed by Luo et al. (2023), adopting a DP prior with base measure $G_{0}$ and concentration parameter $\alpha>0$ allows posterior draws of the data-generating distribution to place mass both on the observed data and on new atoms drawn from $G_{0}$. Consequently, posterior predictive distributions for the marginal dose–response function $\mu(d)$ are no longer confined to the empirical support of the observed doses. This enables uncertainty propagation in regions of the dose space that are sparsely observed or lie near the extremes of the exposure distribution. Moreover, the base measure $G_{0}$ can carry weak structural information and can be interpreted as prior knowledge about the underlying data-generating process. In addition, the parameter $\alpha$ and $G_{0}$ can serve as a regularization mechanism, stabilizing posterior inference in regions with sparse data by shrinking toward the base measure while still allowing the data to dominate where information is abundant.
Therefore, we extend the Bayesian bootstrap to a more general DP specification. If we specify a priori $P\sim DP\left(\alpha,G_{0}\right)$ where $\alpha>0$ is the concentration parameter and $G_{0}$ is the base measure. In light of data $\left(z_{1},\ldots,z_{n}\right)$, the resulting posterior distribution of $P$ is $DP\left(\alpha_{n},G_{n}\right)$, where $\alpha_{n}=\alpha+n$ and $G_{n}(\ldotp)=\alpha G_{0}(\ldotp)\big/{(\alpha+n)}+{\sum_{j=1}^{n}\delta_{z_{j}}\left(\cdot\right)}\big/{(\alpha+n)}$. Specifically, we can generate the collection of data from the DP via $\left\{z_{j}\right\}_{j=1}^{\infty}\sim G_{n}$ and $\left\{p_{k}\right\}_{j=1}^{\infty}\sim StickBreaking(\alpha_{n})$; the standard stick-breaking algorithm (Sethuraman, 1994) generates the weights by a transformation of the collection $\{V_{k}\}_{k=1}^{\infty}$ where $V_{j}\sim Beta(1,\alpha)$ are independent, with $p_{1}=V_{1}$ and for $j=2,3,\ldots$, $p_{j}=V_{j}\prod_{k=1}^{j-1}(1-V_{k})$. Therefore, to obtain a posterior variate, we can sample $\left\{z_{j}^{s}\right\}_{j=1}^{\infty}\sim G_{n}$ and $\left\{p_{j}\right\}_{j=1}^{\infty}\sim StickBreaking(\alpha_{n})$, then solving
$\sum_{j=1}^{\infty}p_{j}\mathbf{U}_{1}\left(x^{s}_{j},d^{s}_{j},\gamma\right)=\mathbf{0},\sum_{j=1}^{\infty}p_{j}\mathbf{U}_{2}\left(z^{s}_{j},\xi,e(x_{j},\widehat{\gamma})\right)=\mathbf{0}.$ (7)
Although this is an infinite sum, the $p_{j}$ decreases in expectation as $j$ increases and eventually becomes numerically negligible. Because of this decay, the infinite sum can be truncated at some finite $J$, such that $p_{J}<\epsilon$ for a small tolerance $\epsilon$. The uncertainty about $\xi$ is properly addressed in this ‘forward’ calculation. We also layout this DP specification for the longitudinal analysis in Algorithm 1. In this formulation, we do not directly specify a prior for $\xi$ but instead it is a functional of $p_{i}$, where we carry out a fully Bayesian inference for $p_{i}$. This approach is computationally more efficient because it relies purely on optimization instead of MCMC, which provides a ‘shortcut’ to fully Bayesian causal inference. The validity and theoretical justification of this approach have been thoroughly examined in cross-sectional settings in Luo et al. (2023).
In this setting, once a new dataset is generated, we first estimate the PS based the GEE related to $\mathbf{U}_{1}$, and then plug this point estimate into the GEE associated with $\mathbf{U}_{2}$. This procedure induces a random empirical measure that weights both the GPS and outcome estimating equations (with the same resampling weights) so that uncertainty is propagated coherently through the two-stage system. Both the GPS model and the outcome model are specified non-parametrically and weighted with the same stick-breaking weight. This distinguishes our approach from the previous work in Saarela et al. (2015); Liu et al. (2020). This is known as a two-step procedure (Kaplan and Chen, 2012) and is commonly used in Bayesian causal inference to avoid model feedback, as PSs should be estimated without incorporating outcome information in order to preserve their balancing properties (Rubin, 2007; McCandless et al., 2010; Saarela et al., 2015). This procedure has been shown to yield preferable estimation performance in practice (Stephens et al., 2023). They further showed that weighting both models with the same Bayesian bootstrap weights, analogous to our approach of using a common stick-breaking weight, yields valid coverage. Our Bayesian implementation in the longitudinal setting performs resampling at the subject level rather than at the observation level. Specifically, we assign a single stick-breaking weight to each study unit, which is held fixed across time. When a study unit is resampled, their entire longitudinal trajectory $\{D_{it},X_{it},Y_{it}\}_{t=1}^{T}$ is included or generated from the base measure in the resampled dataset. This subject-level resampling strategy preserves the within-subject dependence structure across repeated measurements and aligns with the marginal modeling perspective underlying GEE. By contrast, resampling study unit observations would implicitly impose independence across time within subjects and would be incompatible with the longitudinal dependence encoded in the estimating equations.
In terms of GEE-based identification under this framework, Luo et al. (2023) show that, under suitable regularity conditions, the generalized Bayesian bootstrap procedure built upon GEE or more general loss functions yield posterior distributions that concentrate asymptotically around the corresponding frequentist M-estimators. In particular, when the treatment model is correctly specified, the GEE used to define the GPS provides unbiased estimating equations for the target marginal parameters. Consequently, the resulting posterior for the APO is supported on the identified parameter space characterized by Equation (1) and concentrates around the corresponding frequentist estimand, which coincides with the true marginal dose–response function under identification assumptions stated in Section 2.1.
Data: $z_{1:n}=(z_{1},\ldots,z_{n})$
Given $\alpha$ and $G_{0}$, for $s$ to $1:S$ do
[…]
  • Sample $\{p_{k}^{s}\}$ from a stick-breaking process with $\alpha_{n}=\alpha+n$;
  • Estimate the PS based the GEE related to $\mathbf{U}_{1}$ in (7) with the data $\{z^{s}_{k}\}$ and the weight $\{p_{k}^{s}\}$;
  • Solve the GEE related to $\mathbf{U}_{2}$ in (7) and obtain $\xi^{s}$ with the data $\{z^{s}_{k}\}$, the weight $\{p_{k}^{s}\}$ and the estimated PS;
  • Compute the APO, $\zeta^{s}(d)$, at a given dose level $d$ based on the sample $\xi^{s}$ and the estimated PS.
end for
return $(\zeta^{1}(d),\ldots,\zeta^{S}(d))$.
[6pt]
Algorithm 1 Algorithm for generating posterior samples of the APO at a given dose level $d$ based on the generalized Bayesian bootstrap (DP specification).
3 Simulations
In this section, we examine the performance of the methods described in Section 2 in different aspects. The code for the simulation is publicly available on our GitHub page at https://github.com/yumcgill/Bayes_DoseResponse. We consider the following models for our simulation studies.
  • WOR-BB: A weighted outcome approach in which stabilized inverse GPS weights, $w_{t}=f(d)/\widehat{e}_{t}(d,X_{t})$, are incorporated into the GEE for the outcome model. The marginal density of the treatment $f(d)$ is estimated nonparametrically using a kernel density estimator. This formulation reduces variance by centering the weights around one and preventing extreme weights when the GPS is small. Uncertainty in both models is propagated by the Bayesian bootstrap framework, i.e., repeatedly resampling the weight $\mathbf{p}\sim\text{Dirichlet}\left(1,\ldots,1\right)$.
  • COV-BB: A covariate-adjusted outcome approach in which the estimated GPS, $\widehat{e}_{t}(X_{t})$, is included directly as an additional covariate in the outcome model. Posterior uncertainty is propagated using the Bayesian bootstrap framework.
  • WOR-DP: A weighted outcome approach using stabilized inverse GPS weights in the outcome GEE. Uncertainty in both models is propagated by repeatedly resampling from a stick-breaking process with concentration parameter $\alpha=5$ outlined in Section 2.3 and Algorithm 1.
  • COV-DP: A covariate-adjusted outcome approach in which the estimated GPS is included as a covariate in the outcome model. Uncertainty in both models is propagated by repeatedly resampling from a stick-breaking process with concentration parameter $\alpha=5$.
  • Bayes-MSM: A Bayesian MSM in which the causal dose–response function is modeled directly as a parametric function of the dose. Stabilized inverse GPS weights, $w_{t}=f(d)/\widehat{e}_{t}(d,X_{t})$, are incorporated through a weighted likelihood to remove time-dependent confounding. The GPS and outcome models are both modeled parametrically. The outcome model is specified as a cubic spline regression of the outcome on dose, with non-informative priors placed on all regression coefficients and variance parameters. Posterior inference for both GPS and outcome models is obtained via MCMC using Stan (Stan Development Team, 2025). Note this formulation is not a Bayesian analysis based on a fully specified joint likelihood factorization for all observed processes with independent priors on treatment and outcome model parameters (see discussion in Saarela et al., 2015); rather, it follows a Bayesian working-model approach in which weighting is used to target the causal estimand directly.
Specifically, we outline the procedure for generating posterior samples of the APO at a given dose level $d$ using the proposed DP–based method:
  1. 1.
    Resampling under the DP prior. For each Monte Carlo iteration $s=1,\ldots,S$, we first resampled with replacement from the original dataset, obtain the new dataset with the number of study units $N=100$. This step can be viewed as the DP with $\alpha=0$. For each resampled study unit, a random draw determines whether the study unit’s outcome, $y$, is taken directly from the base distribution, with probabilities $\alpha/(\alpha+n)$. For study units drawn from the base distribution, new outcome values are simulated from a normal distribution with mean given by the model-based predicted value and a standard deviation of 0.5. For each study unit, the predicted mean is derived from the outcome model specified in the corresponding simulation example.
  2. 2.
    Weight generation via stick-breaking. Given the resampled data, a corresponding set of weights $\{p_{k}^{(s)}\}$ is drawn from a stick-breaking process with concentration parameter $\alpha_{n}=\alpha+n$. These weights are assigned to the resampled study units and carried forward into the analysis.
  3. 3.
    GPS estimation (first stage). Using the resampled data from Step 1 and weights from Step 2, the GPS is estimated by solving GEE associated with $\mathbf{U}_{1}$. This stage is performed independently of the outcome model in order to avoid feedback from the outcome into the PS estimation.
  4. 4.
    Outcome model estimation (second stage). Conditional on the estimated GPS, the GEE associated with $\mathbf{U}_{2}$ in Equation (2) is solved using the same resampled data and weights, yielding a posterior draw $\xi^{(s)}$ of the structural parameters.
  5. 5.
    Causal estimand computation. Based on the parameter draw $\xi^{(s)}$ and the estimated GPS, we can predict outcome, $\zeta^{(s)}(d)$, at specific dose level $d$. Repeating the above steps for $s=1,\ldots,S$ yields posterior samples
    $\{\zeta^{(1)}(d),\ldots,\zeta^{(S)}(d)\},$
    which are used to summarize uncertainty in the APO through posterior means, credible intervals, and coverage rates.
3.1 Example 1
In this example, we conduct a simulation study on a sample size of 100 study units, each with 10 time points and a total of 1000 observations. We generate the data via the following relationships:
$X_{1it}\sim\mathcal{N}\left(0.2,0.1\right),X_{2it}$ $\sim\mathcal{N}\left(1,0.6\right),U_{i}\sim\mathcal{N}\left(0.2,0.1\right)$
$D_{it}|X_{1it},X_{2i},U_{i}$ $\sim\mathcal{N}(5+4X_{1it}+2X_{2i}+U_{i},1)$
$Y_{it}|D_{it},X_{1it},X_{2i},U_{i}$ $\sim\mathcal{N}\left(20\exp\left[D_{it}+X_{1it}-0.25X_{2it}+0.5U_{i}\right],1\right)$
$D$ is a continuous treatment variable, and $X_{1}$, $X_{2}$ are time-varying confounders and $U$ is a time-invariant confounder for $D$ with a nonlinear relationship in $Y$. The goal of this causal setting is to estimate the APO, i.e., the outcome at a specific dose level. For the GEE associated with the first stage of the GPS estimation, $\mathbf{U}_{1}$, we include only the covariates $X_{1it}$ and $X_{2it}$. For the GEE associated with the outcome model, $\mathbf{U}_{2}$, we use the outcome in log scale, $\log Y_{i}$. Specifically, in the covariate-adjusted (COV) method, we include $D_{it}$ and a cubic spline of the estimated GPS, allowing for a flexible, nonlinear relationship between the GPS and the outcome. In the weighted outcome regression (WOR) method, we include a cubic spline of $D_{it}$ only. For the DP-based method, the predicted mean is obtained from a GEE outcome model that includes $D_{it}$ and a cubic spline of the estimated GPS.
We first simulate one dataset and generate 1000 posterior samples. Figure 2 shows the 1000 posterior dose-response curves with observed data points. All methods recover the general monotonic increasing relationship between dose and outcome, but there are differences in tail behavior and uncertainty quantification. The WOR-based approaches exhibit some deviation from the true dose–response curve, particularly in the high dose region. However, these regions are characterized by larger variance, leading to wider credible intervals that maintain coverage of the true curve. The DP-WOR method yields slightly wider intervals than BB-WOR, reflecting the additional uncertainty introduced by the more flexible prior on the data-generating distribution. In contrast, the COV-based approaches produce smoother dose–response estimates with relatively narrow credible intervals and remain close to the true curve even in regions with limited data support. This suggests that the COV-based formulation provides a more stable and reliable estimation framework, particularly in settings with small sample sizes. Finally, the Bayes-MSM approach also yields smooth estimates with comparatively tight credible intervals, but shows noticeable departures from the true curve in both tails, highlighting its sensitivity to parametric assumptions when extrapolating beyond well-supported regions of the data.
Subsequently, we repeat this simulation on 1000 datasets, each consisting of 1000 entries, and evaluate the performance on three given doses. The results are shown in Table 1, with averages of the posterior means (Est), standard deviations (SD) and coverage rates based on with the 2.5% and 97.5% posterior sample quantiles. The true APO at dose $d$ was estimated by Monte Carlo integration with 100,000 observations, corresponding to 10,000 study units with 10 measurements each. The APO estimates are calculated for doses 6, 8, and 10. The APO estimates in all models are represented via a logarithmic transformation, and the results are also presented in the logarithmic scale. All methods yield the average of the posterior means close to the true values. The COV-based approaches are more efficient than the WOR-based ones, with smaller standard deviations. The DP-based methods attain near-nominal coverage across all doses considered, whereas COV-BB achieves nominal coverage only at doses with adequate data support and exhibits undercoverage at sparsely observed doses. The WOR-BB method shows consistently lower coverage across doses, while the Bayes-MSM displays undercoverage at doses with limited data support. Overall, this example indicates that the proposed DP-based methods are well suited for longitudinal data with repeated measures, providing valid uncertainty quantification even in low-data regions.
Figure 2: Example 1: Posterior predictive dose–response curves for the Gaussian outcome estimated using the Bayesian bootstrap, Dirichlet process, and Bayesian parametric MSM approaches.
Table 1: Simulation results for the dose–response function at three dose levels, reporting the average posterior mean (Est) with associated standard deviation (SD), and empirical coverage rate (Coverage, %). Results are summarized over 1,000 simulation replicates.
Example 1 (Results are presented in the logarithmic scale)
Method dose = 6 dose = 8 dose = 10
(9.046) (11.046) (13.046)
COV-$B$ Est 8.987 11.045 13.104
SD 0.046 0.036 0.048
Coverage 75.4 94.4 75.7
WOR-$B$ Est 8.959 11.037 13.118
SD 0.217 0.145 0.205
Coverage 77.6 84.3 78.3
COV-DP Est 8.987 11.045 13.103
SD 0.086 0.055 0.088
Coverage 98.2 98.7 97.4
WOR-DP Est 8.964 11.039 13.115
SD 0.188 0.128 0.176
Coverage 96.9 98.8 96.7
Bayes-MSM Est 8.994 11.043 13.097
SD 0.091 0.068 0.096
Coverage 84.5 92.3 82.2
Example 2
Method dose = 6 dose = 8 dose = 10
(9.212) (13.743) (20.503)
COV-$B$ Est 9.175 13.737 20.570
SD 0.128 0.131 0.195
Coverage 93.7 93.9 91.8
WOR-$B$ Est 9.097 13.739 20.686
SD 0.547 0.498 0.972
Coverage 82.8 85.5 81.8
COV-DP Est 8.879 13.240 19.817
SD 0.202 0.182 0.381
Coverage 98.2 91.5 98.0
WOR-DP Est 8.763 13.220 19.879
SD 0.469 0.454 0.788
Coverage 94.9 94.2 97.1
Bayes-MSM Est 9.103 13.686 20.494
SD 0.286 0.261 0.446
Coverage 81.2 91.1 90.4
3.2 Example 2
In this example, we demonstrate inference for simulated outcome as the count data, similar to our real data. All the variables simulated from the same distribution as in Example 1, except the outcome which is generated from a Poisson distribution:
$Y_{it}|D_{it},X_{1it},X_{2i},U_{i}\sim\text{Poisson}\left(\exp\left[1+0.2D_{it}+0.005\frac{X_{1it}}{100}-0.002\frac{X_{2it}}{100}+0.1U_{i}\right]\right).$
Similar to the first example, for the GEE associated with the first stage of GPS estimation, we include only the covariates $X_{1it}$ and $X_{2it}$. For the outcome model, we fit a Poisson GEE for $Y_{i}$. In the COV method, the model includes $D_{1it}$ and a cubic spline of the estimated GPS. In the WOR method, the outcome mean is modeled using a cubic spline of $D_{1it}$, with the estimated GPS incorporated as a weight. For the DP-based method, the predicted mean is obtained from a GEE outcome model including $D_{it}$ and a cubic spline of the estimated GPS.
Figure 3 demonstrates the 1000 posterior dose-response curves with observed data points. Both BB-based and DP-based approaches closely align with the true dose–response curve, with only small deviations at the high-dose tail. The WOR-based formulations exhibit wider 95% credible intervals, reflecting increased uncertainty in regions with limited data support. Similar with the previous example, the DP-based methods produce slightly wider intervals at extreme doses, capturing additional uncertainty induced by the more flexible prior. The Bayes-MSM method shows the most deviation of the posterior mean from the true curve at high doses, suggesting that MSM can lead to certain bias in regions where data are sparse.
Table 1 summarizes the results of 1,000 simulation runs at three specified dose levels. All methods both give roughly unbiased estimates for all dose levels. WOR-based methods again exhibit larger standard deviations, with WOR-DP achieving near-nominal coverage, while WOR-BB shows systematic undercoverage across all doses. The COV-based methods maintain good coverage at all dose levels. Similar to the previous example, Bayes-MSM shows undercoverage at the dose extremes, highlighting its sensitivity in regions with limited data support as we observed in Figure 3.
Figure 3: Example 2: Posterior predictive dose–response curves for the Poisson outcome estimated using the Bayesian bootstrap, Dirichlet process, and Bayesian parametric MSM approaches.
4 Application: Quantifying the causal impact of mass transit on COVID-19 transmission
As discussed in the introduction, empirical studies quantifying the causal impact of mass public transport on COVID-19 transmission are scarce. For estimation, we have monthly (longitudinal) city level data available which measure Covid cases, mass transit ridership, mobility containment measures, and a set of mobility characteristics for the period between 1 March 2020 and 23 January 2022. Longitudinal data provide the temporal resolution needed to understand how public transit ridership and COVID-19 incidence evolve and intersect. Most use aggregate measures of mobility from mobile phone data sources such as those that were made publicly available by Google Mobility and Apple Maps during the pandemic, and issues of causality have received little attention. In our review, we have identified just one empirical study by Romero García et al. (2023) which uses causal analysis methods to assess the impact of travel specifically on metro networks (referred to as subways in the paper) on the transmission of COVID-19 in Valencia, Spain. In the following sections, we provide a summary of empirical studies undertaken on the impact of mobility on COVID-19 transmission, beginning with an overview of literature using non-causal analysis methods, followed by a summary of the limited literature adopting causal methods.
4.1 Non-causal empirical studies
There is a vast number of empirical studies that use non-causal statistical analysis methods to study the association of mobility with COVID-19 transmission as measured by either the number of COVID-19 cases or the virus reproduction number. Generally, most studies conclude that an increase in mobility is positively associated with an increase in COVID-19 transmission, and restrictions on mobility are associated with reductions in COVID-19 transmission. Studies using statistical correlation methods are prevalent in the literature. For example, Jia et al. (2020) reported a positive linear correlation between mobility and case incidence in China at the early stages of the pandemic. Badr et al. (2020) analyzed mobility patterns in the USA and also reported positive linear correlations between mobility and COVID-19 case rates. Kissler et al. (2020) used Pearson’s linear correlation coefficient and reported that their estimates of COVID-19 prevalence were negatively associated with reductions in mobility in their analysis of New York City.
Studies adopting different forms of multivariate regression methods are also prevalent, and here we present a selection of examples. Li et al. (2021) used multivariate linear regression and reported that the reproduction number of the virus increased with higher levels of mobility at retail and recreation places, workplaces, and transit stations in the UK. Kraemer et al. (2020) applied a generalized linear regression model framework and found that mobility was positively correlated with case incidence in China prior to travel restrictions, but the relationship was no longer statistically significant after the implementation of restrictions. McGrail et al. (2020) undertook a cross-country analysis of 134 countries and 50 US states. The authors used a generalized linear mixed effects model and reported that mobility and infection rates reduced after the imposition of social distancing restrictions. Islam et al. (2020) also performed a cross-country analysis of 149 countries using a two-stage linear regression analysis. It was reported that case incidence reduced with the imposition of public transport closures, school closures, work closures, restrictions on mass gatherings, and blanket lockdowns. Xiong et al. (2020) reported a positive relationship between mobility and COVID-19 cases in the US using structural equation modeling with dynamic panel and time-varying effects at the level of US counties. Manzira et al. (2022) used a multivariate time series regression model with autoregressive lag effects and reported a positive association between mobility and the incidence of COVID-19 in the city center areas of Dublin, Ireland.
4.2 Causal empirical studies
There are a limited number of studies that adopt causal analysis methods to quantify the impact of mobility on COVID-19 case transmission. Romero García et al. (2023) undertook the only known analysis of the impact of metro travel on COVID-19 case rates. Romero García et al. (2023) performed a small-scale analysis of the impact of the locations of subway stations and hospitals on COVID-19 cases in Valencia, Spain. They first partitioned the region using Voronoi diagrams using subway locations and hospitals as segmentation variables, and then applied the Granger causality test. A positive association was reported between interchange subway stations and COVID-19 cases, while transmission from hospital locations generated more mixed results.
Steiger et al. (2021), Nugent and Balzer (2023), Cho et al. (2022), and Chernozhukov et al. (2021) used aggregate indicators of mobility from mobile phone data records unattributed to specific travel modes to assess causality between mobility and COVID-19 cases. Steiger et al. (2021) estimated the causal impact of factors affecting the number of new cases of COVID-19 in Germany using a DAG and a negative binomial regression model. The authors found a positive causal impact between mobility in retail and recreational spaces and workplaces and COVID-19 cases, and a negative causal impact for essential shopping and mobility within residential areas. Other influential factors included weather, socio-demographic characteristics, and public awareness of COVID-19. Nugent and Balzer (2023) adopted the causal modified treatment policy method with the Super Learner machine learning algorithm to quantify the impact of mobility on COVID-19 cases in US counties. The authors found that unadjusted estimates suggested that decreasing mobility is causally linked to decreasing COVID-19 cases, but found that their adjusted estimates gave mixed results indicating potential confounding in the unadjusted estimates, which were most likely attributed to unknown differences in county characteristics. Cho et al. (2022) also analyzed the causal relationship between COVID-19 cases and mobility in the US. They adopted an ensemble empirical mode decomposition method with causal decomposition and reported that mobility levels have a causal relationship with long-term variations in COVID-19 cases rather short-term variations. Chernozhukov et al. (2021) used a causal structural model framework to test a range of counterfactual experiments including mobility changes and their impact on COVID-19 cases in the US. The authors find that reductions in mobility due to stay at home orders and business closures were effective in reducing case numbers, while the impacts of reductions in mobility due to school closures were inconclusive.
4.3 Data
The Community of Metros (COMET) benchmarking group administered by the Transport Strategy Centre at Imperial College London have provided ridership data from 8 international member metro networks. In keeping with commercial confidentiality requirements, the metro identities have been anonymized and are referred to with a numerical index ranging from 1 to 8 throughout this paper. hlTherefore, the study units correspond to eight distinct geographical locations. The ridership data reports the numerical ridership figures for each day. The data are recorded from before the onset of the COVID-19 pandemic to early 2022, with the earliest and latest dates in the data set being 1 March 2020 and 23 January 2022.
Data on COVID-19 case numbers correspond to the city that each metro is located in and refer to daily case numbers. The data on case numbers have been obtained from the Oxford University COVID-19 Government Response Tracker (OxCGRT) (Hale et al., 2021). Additional data from the OxCGRT database included in the analysis are the number of COVID-19 deaths per day, and a metric termed the Stringency Index which represents the severity of policy control measures implemented by governments during COVID-19 to limit transmission. The Stringency Index is a numeric index ranging from 0-100 with higher values corresponding to stricter control measures. We also include data on vaccinations which represent the cumulative total of vaccinations delivered on a daily basis from the Our World in Data database (Mathieu et al., 2021). Lastly, we include a series of variables representing all forms of mobility from Google Mobility (Google LLC, 2022), specifically: residential, workplace, retail and recreation, grocery and pharmacy, and parks.
For context, a summary of all variables including their descriptive statistics aggregated across all metros in the analysis is given in Table 2. The ridership is transformed in a logarithmic scale, and the number of cases are transformed with $\log(x+1)$ scale. From our preliminary analysis, the distribution of ridership is highly skewed and therefore we take the right $\log$ scale. After the transformation, the distribution looks more normal, and the ridership in $\log$ scale will be the treatment variable. For the outcome variable, we will consider to use cases without transformation and in $\log(x+1)$ scale.
Table 2: Descriptive statistics of metro ridership data
Variable Description Min. Max. Mean Std. Dev.
Ridership (loglog scale) Number of passengers per day in log scale 10.17 15.15 13.02 1.07
Cases ($log(x+1)$ scale) Number of confirmed COVID-19 cases in $log(x+1)$ scale 4.52 14.31 10.24 1.81
Deaths Number of confirmed COVID-19 deaths 12 121201 40167 37874.10
Stringency index Stringency index ranging from 0-100 from OxCGRT. Represents degree of severity of policy measures to contain COVID-19 transmission. 20.37 98.46 62.44 16.04
Total vaccinations per hundred Number of total vaccinations i.e. sum of all doses per hundred people of population. For vaccines that require multiple doses, each individual dose is counted. 0 238 56.40 71.87
Retail and recreation Google mobility data for mobility changes relative to a baseline regular (pre-Covid) day for retail and recreation places -97 36 -37 22.33
Grocery and pharmacy Google mobility data for mobility changes relative to a baseline regular (pre-Covid) day for grocery and pharmacy places -91 99 -9 20.16
Parks Google mobility data for mobility changes relative to a baseline regular (pre-Covid) day for parks -98 294 -3 51.67
Transit stations Google mobility data for mobility changes relative to a baseline regular (pre-Covid) day for transit stations -93 61 -41 21.80
Workplaces Google mobility data for mobility changes relative to a baseline regular (pre-Covid) day for workplaces -92 31 -35 21.77
Residential Google mobility data for mobility changes relative to a baseline regular (pre-Covid) day for residential places -9 48 13 8.75
4.4 Results
To estimate the dose-response relationship between the ridership and COVID-19 cases, we employed the proposed Bayesian method to analyze two different outcomes: the raw case counts, modeled using a log link with the Poisson GEE, and the log-transformed case counts, $log(x+1)$, modeled under an identity link with the Gaussian GEE. For both outcomes, we specified an independent working correlation structure and included all relevant confounders along with a cubic spline function of the GPS as covariates. Similar to the simulation, for the GPS model, we used a GEE model that adjusts for all time-varying confounders. We generated a total of 5,000 posterior samples. For the DP method, we generate a synthetic dataset of size $N=200$ using a procedure similar to that employed in the simulation studies. Due to the relatively small sample size, the stick-breaking weights in the DP decreased rapidly, falling below $10^{-8}$ after approximately 200 iterations. Figure 4 presents posterior predictive dose–response curves estimated using COV and WOR methods. Across all specifications, the estimated dose–response relationship indicates an overall positive and monotone causal effect of mobility on COVID-19 cases, with higher ridership associated with higher expected case counts. Under the COV-based approach, the estimated dose–response function increases for both log-transformed and count outcomes, although uncertainty is larger in certain dose ranges (notably between 10–12 and 14–15) due to reduced data density. The DP-based approach shows a more stable curve as the synthetic dataset have a larger sample. When using case counts as the outcome, the COV-based method shows a generally exponential increase in cases as the logarithm of ridership rises. However, the DP-based methods appear to overcompensate in certain regions, resulting in some non-smooth curves likely because some portion of the resampled data is drawn from areas that deviate from the main trend. In contrast, the WOR estimator shows a rapid increase in cases at lower dose levels followed by a plateau and slight decline at higher doses, indicating sensitivity to reweighting in sparsely populated regions. In regions with sparse data, the DP-based approach produces more stable estimates, suggesting improved performance in low-density areas. In summary, the number of cases tends to increase as the dose level, representing that public transport usage rises, but this relationship is not strictly linear. While both the log-transformed and count-based outcome suggest a strong upward trend in case numbers at lower dose regions, the log-transformed model shows a more moderate increase due to its smoothing effect, whereas the count model captures a steeper, non-smooth and exponential rise.
Figure 4: Application: Posterior predictive for dose–response estimates for ridership ($\log$ scale) and COVID-19 cases. Panels labeled $\log(x+1)$ use Gaussian GEE models for log-transformed cases, while panels labeled $\log$ scale use Poisson GEE models with posterior predictive samples rescaled to the $\log$ scale. Results are shown for Bayesian bootstrap and DP inference under covariate-adjusted and weighted outcome regression specifications.
5 Discussion
In this paper, we introduced a non-parametric Bayesian framework for estimating longitudinal causal dose-response relationships in the presence of repeated measures. The proposed framework addresses Bayesian inference under model misspecification by adopting a GEE perspective. Instead of specifying a fully parametric likelihood, the target estimand is defined through moment conditions that characterize the marginal causal parameters of interest, and uncertainty is quantified by placing a nonparametric prior on the underlying data-generating distribution. Specifically, this approach extends traditional Bayesian methodology by integrating the generalized Bayesian bootstrap within a GEE framework, thereby explicitly accounting for the temporal structure of longitudinal data. Both the GPS and the outcome mean are modeled using GEEs, and the GPS is incorporated into the outcome model in two ways: as an additional covariate or through stabilized inverse weighting. Simulation results indicate that both Bayesian dose–response estimators are approximately unbiased and achieve near-nominal posterior coverage across a range of scenarios when using the DP specification. However, the weighting-based estimator exhibits larger posterior variance, a behavior consistent with other inverse-weighting approaches for continuous treatments, where exposure binning is often required to mitigate variability (Naimi et al., 2014). In our empirical study, we investigated the impact of metro ridership on COVID-19 case counts across major urban centers during the pandemic. The results suggest that increased metro ridership contributes to the spread of COVID-19 up to a certain exposure threshold, beyond which the effect plateaus. These findings have important implications for pandemic-related urban policy, particularly in informing lockdown strategies and optimizing transportation systems in future public health crises.
One appealing aspect of Bayesian estimators is their flexibility, as they allow for flexible approaches with minimal assumptions about the dose-response relationship, and for the prior knowledge to be added in the model. The proposed method can easily handle conditionally censored data by incorporating time-varying censoring weights. Bayesian methods are increasingly influential in applied causal inference, offering the ability to make direct probabilistic statements about treatment effects and to assess sensitivity with respect to alternative priors or expert-driven inputs. Despite these advantages, Bayesian causal methods for longitudinal data remain underdeveloped (Liu et al., 2020). In this paper, we proposed a scalable Bayesian approach that effectively handles longitudinal data and is easily implementable. Our proposed Bayesian dose-response estimation is derived under the SUTVA assumption, assuming no interference. In our motivating COVID study, this assumption translates to each unit’s monthly metro-ridership exposure being assumed to affect only its own subsequent COVID-19 risk, which can be violated in application. Methodological extensions that relax SUTVA are emerging (O’Riordan and Gilligan-Lee, 2025). Future work will adapt these ideas to longitudinal, continuous-dose settings, enabling joint estimation of direct and indirect (spillover) ridership effects on COVID-19 dynamics. Despite this limitation, the proposed estimator fills an important methodological gap by coupling non-parametric Bayesian inference with repeated-measures data and a time-varying continuous exposure. It remains fully appropriate for applications where interference is unlikely (e.g., individual risk-factor studies in chronic-disease epidemiology). Furthermore, our framework offers a robust foundation that broader spillover-aware extensions can build upon.
Acknowledgement
Yu Luo was supported by the Engineering & Physical Sciences Research Council (EPSRC) Grant EP/Y029755/1. Kuan Liu was supported by the Canadian Institutes of Health Research (CIHR) Grant AD7200183.
Appendix A GPS estimation with an unobserved, time-invariant confounder
In GPS estimation, we model the conditional mean of treatment given observed covariates. Let $D_{it}$ denote the treatment at time $t$ for study unit $i$, $X_{it}$ the observed time-varying confounders, and $U_{i}$ an unobserved, time-invariant subject-specific effect. The true treatment assignment model is
$D_{it}=\gamma X_{it}+U_{i}+\varepsilon_{it},\quad\mathbb{E}(\varepsilon_{it}\mid X_{it},U_{i})=0,$ (8)
with potentially correlated repeated measurements within subjects. GEE estimation proceeds by specifying a marginal mean model
$\mu_{it}(\gamma)=\mathbb{E}(D_{it}\mid X_{it})=\gamma X_{it}.$ (9)
The key property of GEE is that it targets the marginal mean directly, rather than the conditional mean given $U_{i}$. By the law of iterated expectations, the marginal mean is
$\mathbb{E}(D_{it}\mid X_{it})$ $=\mathbb{E}_{U_{i}|X_{it}}\big[\mathbb{E}(D_{it}\mid X_{it},U_{i})\mid X_{it}\big]$
$=\mathbb{E}_{U_{i}|X_{it}}[\gamma X_{it}+U_{i}\mid X_{it}]$
$=\gamma X_{it}+\mathbb{E}(U_{i}\mid X_{it}).$ (10)
Under the standard GEE assumption that $U_{i}$ is mean-independent of the observed covariates,
$\mathbb{E}(U_{i}\mid X_{it})=0,$ (11)
the marginal mean reduces to
$\mathbb{E}(D_{it}\mid X_{it})=\gamma X_{it},$
which matches the model in Equation (9). Consequently, $U_{i}$ does not appear in the marginal mean and does not induce bias in the GEE estimator of $\gamma$. The GEE estimator $\widehat{\gamma}$ solves an estimating equation whose expectation equals to 0 at true parameter $\gamma$. Therefore, by standard GEE theory, $\widehat{\gamma}\overset{p}{\to}\gamma$ given the correct specification of the marginal model regardless of the choice of the working correlation.
In summary, unobserved time-invariant subject effects $U_{i}$ do not bias the marginal GPS estimate as long as $\mathbb{E}(U_{i}\mid X_{it})=0$, because they are averaged out in the marginal mean. Therefore, consistent estimation of the GPS is guaranteed under correct specification of the marginal mean model, independent of the correlation structure or the presence of $U_{i}$.
Appendix B Estimation of regime-specific average potential outcomes in longitudinal settings
B.1 Longitudinal regime setup and notation
In causal inference for the longitudinal setting, for $t=0,\ldots,T$, let $D_{t}\in\mathbb{R}$ denote the (possibly continuous) dose at time $t$, $\bar{D}_{t}=(D_{t-c},\ldots,D_{t})$ the treatment dosing history where $1\leq c\leq t$ specifies the length of the relevant treatment history. Therefore, $\bar{D}_{t}$ represents the sequence of recent doses up to and including time $t$, which may influence the potential outcome $Y_{t}$. Let $\bar{X}_{t}=(X_{0},\ldots,X_{t})$ denote the covariate history and $X_{t}$ may include measured lagged summaries (e.g., $Y_{t-1}$ recent policy indices, or recent mobility indicators), so that $X_{t}$ functions as a sufficient “state” capturing the relevant observed history prior to assigning $D_{t}$. Let $U$ denote an unobserved time-invariant baseline covariate that may affect both treatment and outcomes. Using the potential outcome framework in the longitudinal setting, we set $Y_{t}(\bar{d}_{t})$ to denote the counterfactual outcome at time $t$ that would be observed if a study unit were assigned the dose history $\bar{d}_{t}=(d_{1},\ldots,d_{t})$ up to time $t$. The primary focus of our inference is on APOs for a population under a hypothetical intervention, given the dose history $\bar{d}_{T}=(d_{1},\ldots,d_{T})$ over $T$ intervals. That is, the estimand of interest is $\mathbb{E}\!\left[Y_{t}(\bar{d}_{t})\right]$.
To identify the APO, we require several assumptions. First, we assume that there is no interference across units, i.e., for each study unit $i$, $Y_{t,i}(\bar{d}_{1},\ldots,\bar{d}_{n})=Y_{t,i}(\bar{d}_{i})$, corresponding to the SUTVA. Second, we assume consistency, such that if a study unit follows dose history $\bar{D}_{t}=\bar{d}_{t}$, then the observed outcome satisfies $Y_{t}=Y_{t}(\bar{d}_{t})$ for all $t=1,\ldots,T$. Third, we assume sequential ignorability: at each time $t=1,\ldots,T$, conditional on $\bar{X}_{t}$ and past doses $\bar{D}_{t-1}$, the treatment assignment at time $t$ is independent of future potential outcomes, $Y_{s}(\bar{d}_{s})\;\perp\!\!\!\perp\;D_{t}\mid\bar{X}_{t},\bar{D}_{t-1},\forall s\geq t$. Fourth, we assume positivity, such that for any $(\bar{x}_{t},\bar{d}_{t-1})$ with positive density, $f(d_{t}\mid\bar{X}_{t}=\bar{x}_{t},\bar{D}_{t-1}=\bar{d}_{t-1})>0$ over the range of doses considered. Finally, we assume that unobserved time invariant effects $U$ and observed time-varying confounders $\bar{X}_{t}$ are uncorrelated, i.e., $\mathbb{E}(U|\bar{X}_{t})=0$. This ensures that, after conditioning on the full observed history, $U$ does not induce bias in either the treatment model $f(d_{t}\mid\bar{X}_{t},\bar{D}_{t-1})$, or the outcome regression or GEE-based estimating equations. Figure 5 depicts a corresponding longitudinal causal structure, allowing for treatment–confounder feedback, lagged covariate effects, and unobserved time-invariant heterogeneity.
Figure 5: Directed acyclic graph for a general longitudinal data-generating process with time-varying treatment $D_{t}$, time-varying confounders $X_{t}$, outcomes $Y_{t}$, and unobserved time-invariant heterogeneity $U$
$X_{t}$$D_{t}$$Y_{t}$$X_{t+1}$$D_{t+1}$$Y_{t+1}$$U$
Following the standard marginal effect parameterization for time-varying treatments and repeated outcomes, the APO at time $t$ is modeled as
$g\!\left(\mathbb{E}\!\left[Y_{t}(\bar{d}_{t})\right]\right)=m(\bar{d}_{t},\bar{x}_{t},u;\,\xi),$
where $g(\cdot)$ is a known link function and $m(\cdot;\xi)$ is a conditional mean function indexed by parameters $\xi$. Marginal structural models (MSMs) estimate such marginal effects by reweighting observations using inverse probability weights to remove time-dependent confounding. We can also incorporate the GPS directly into the outcome model to adjust for time-varying confounding. This approach yields consistent estimates of the APO provided that the PS model is correctly specified. Accordingly, this model can be rewritten as
$g\!\left(\mathbb{E}[Y_{t}\mid\bar{D}_{t},\bar{X}_{t},\widehat{e}_{t}(X_{t})]\right)=m(\bar{D}_{t},\bar{X}_{t};\xi)+h\!\left(\widehat{e}_{t}(\bar{X}_{t},\bar{D}_{t-1})\right),$
where $\widehat{e}_{t}(\cdot)$ denotes the estimated GPS at time $t$ and $h(\cdot)$ is a prespecified adjustment function, which may be estimated using flexible methods such as splines. For continuous treatments, the GPS at time $t$ is defined as the conditional density of the observed dose given the observed history,
$e_{t}(\bar{X}_{t},\bar{D}_{t-1})=f(D_{t}=d_{t}|\bar{X}_{t},\bar{D}_{t-1}).$
Both the GPS and the outcome model can be estimated via GEEs, and their corresponding estimating equations are embedded within the proposed Bayesian framework to enable posterior inference without full likelihood specification.
B.2 Identification of regime-specific APO
Given the assumptions for the regime-specific APO in Appendix B.1, $\mu_{t}(\bar{d}_{t})\equiv\mathbb{E}\!\left\{Y_{t}(\bar{d}_{t})\right\}$ is identified as following.
First, by iterated expectations,
$\mu_{t}(\bar{d}_{t})=\int\mathbb{E}\!\left\{Y_{t}(\bar{d}_{t})\mid\bar{X}_{t}=\bar{x}_{t}\right\}\,f_{\bar{d}_{t}}(\bar{x}_{t})\,d\bar{x}_{t},$
where $f_{\bar{d}_{t}}(\bar{x}_{t})$ is the density of $\bar{X}_{t}$ under the hypothetical intervention that sets $\bar{D}_{t}=\bar{d}_{t}$. Applying sequential ignorability recursively and using consistency on the event $\{\bar{D}_{t}=\bar{d}_{t}\}$ yields
$\mathbb{E}\!\left\{Y_{t}(\bar{d}_{t})\mid\bar{X}_{t}=\bar{x}_{t}\right\}=\mathbb{E}\!\left\{Y_{t}\mid\bar{D}_{t}=\bar{d}_{t},\bar{X}_{t}=\bar{x}_{t}\right\}.$
Under the intervention $\bar{D}_{t}=\bar{d}_{t}$, the covariate process evolves according to the same conditional transition laws as in the observed data, with past doses set to the specified values. Hence
$f_{\bar{d}_{t}}(\bar{x}_{t})=\prod_{s=0}^{t}f\!\left(x_{s}\mid\bar{X}_{s-1}=\bar{x}_{s-1},\,\bar{D}_{s-1}=\bar{d}_{s-1}\right).$
yielding
$\mu_{t}(\bar{d}_{t})=\int\mathbb{E}\!\left\{Y_{t}\mid\bar{D}_{t}=\bar{d}_{t},\bar{X}_{t}=\bar{x}_{t}\right\}\prod_{s=0}^{t}f\!\left(x_{s}\mid\bar{X}_{s-1}=\bar{x}_{s-1},\,\bar{D}_{s-1}=\bar{d}_{s-1}\right)\,d\bar{x}_{t}.$